OpenStudy (anonymous):

What does xy-5=k help me plzz

OpenStudy (blurbendy):

which value are you trying to solve for?

OpenStudy (blurbendy):

variable*

OpenStudy (anonymous):

I'm trying to find for x

OpenStudy (blurbendy):

so you need to get x by itself on the left side of the equation. start by adding five to both sides. xy = k + 5 now you just have to divide both sides by y x = (k + 5) / (y) or x = k/y + 5/y